### v1

# Why In-Context Learning Transformers are Tabular Data Classifiers

Abstract
The recently introduced TabPFN pretrains an In-Context Learning (ICL) transformer on synthetic data to perform tabular data classification. As synthetic data does not share features or labels with real-world data, the underlying mechanism that contributes to the success of this method remains unclear. This study provides an explanation by demonstrating that ICL-transformers acquire the ability to create complex decision boundaries during pretraining. To validate our claim, we develop a novel forest dataset generator which creates datasets that are unrealistic, but have complex decision boundaries. Our experiments confirm the effectiveness of ICL-transformers pretrained on this data. Furthermore, we create TabForestPFN, the ICL-transformer pretrained on both the original TabPFN synthetic dataset generator and our forest dataset generator. By fine-tuning this model, we reach the current state-of-the-art on tabular data classification. Code is available at https://github.com/FelixdenBreejen/TabForestPFN.
1 Introduction
Tabular data classification is widespread across all industries, leading to an increased interest in the research field of deep learning for tabular data [28, 52, 25, 35]. This type of classification involves classifying a target variable based on a set of attributes, which is commonly stored in tabular format. Examples of tabular classification include predicting the existence of chronic kidney disease based on blood test results [34], estimating the click-through rate of advertisements [38], and predicting the stability of pillars in hard rock mines [29]. Despite the significance of tabular data, major breakthroughs in AI as shown on vision [16, 19, 9] and language [48, 3] data have yet to reach the tabular domain. In fact, neural networks are currently outperformed by tree-based machine learning algorithms such as XGBoost [6] and CatBoost [37] in tabular classification tasks [12, 15, 32].
In an attempt to bridge this performance gap, a recent method called tabular prior-data fitted networks (TabPFN) [20] introduces an in-context learning (ICL) [8] scheme, demonstrating promising results [15]. This tabular ICL-transformer can predict test observations with only one forward pass using training observations included in the context. Hollmann et al. [20] generate their pretraining data synthetically, with a focus on creating realistic datasets that act as a “prior”. They make their datasets realistic by carefully crafting correlations between features, introducing variety in the feature importance, and by leveraging structural causal models, which can simulate causal relationships.
Even with realistic datasets, it is questionable why pretraining ICL-transformers would improve classification performance on real-world tabular benchmarks. While pretraining in vision and language has clear purposes by transferring texture and grammar knowledge, explanations from a transfer learning point of view in the field of tabular data do not make sense, as datasets do not share features or labels.
In this work, we claim that the effectiveness of ICL-transformers stems from their ability to form complex decision boundaries. See Figure 1 for an example of such a decision boundary. Intuitively, the complexity is given by how far the decision boundary differs from a simple linear line. Neural networks that train from scratch on tabular data often have overly simple decision boundaries, a phenomenon known as simplicity bias [41], while tree-based methods do not suffer from this [15]. We show that ICL-transformers, in contrast, learn how to create these complex decision boundaries during pretraining. To support our claim, we present three arguments:
  • ICL-transformers trained on a highly complex but unrealistic dataset generator achieve performance similar to TabPFN. We create this novel forest dataset generator based on decision trees, and we call the ICL-transformer pretrained on this data TabForest. Although the generator outputs orthogonal decision boundaries and independent and identically distributed features, the performance of TabForest on individual benchmark datasets is strikingly similar to TabPFN.
  • Pretraining dataset complexity is highly correlated with performance. To show this, we vary the parameters of our forest dataset generator that influence the complexity of the decision boundaries. We observe that higher complexity increases the performance.
  • ICL-transformers can create decision boundaries with high complexity. This property emerges when fine-tuning on a specific dataset, as illustrated in Figure 1. Furthermore, we observe that complexity increases with the size of the model.
Figure 1: Comparison of decision boundaries for the Electricity dataset (OpenML ID 44156). Axis represent features, colors are predicted class probabilities, and dots are test observations. Fine-tuned variants show a higher complexity score (see section 5.4) than non-fine-tuned variants (Zero-shot).
In addition to explaining why ICL-transformers work, we push their performance to new heights. TabForestPFN, the ICL-transformer trained on both the TabPFN and the forest dataset, reaches state of the art on two benchmarks [15, 32]. This ensemble boosts performance by seemingly selecting the best method between TabPFN and TabForest, leveraging the small unexplained differences in performance across various real-world datasets.
We achieve a second major push in performance by switching from zero-shot to fine-tuning. The additional fine-tuning phase improves performance on average, even when the number of observations is less than a thousand. Moreover, both fine-tuning and zero-shot performance highly depends on the context size, where larger contexts yield better results, even far beyond the context size pretrained on. These findings suggest that fine-tuning, using the largest possible context size, should be the default choice when working with ICL-transformers.
In conclusion, ICL-transformers are highly effective tabular data classifiers, capable of creating complex decision boundaries. This new insight advances our understanding of tabular ICL-transformers and opens up new avenues for further research to enhance their performance. With further developments, we anticipate a significant shift in the field of tabular data, moving from tree-based methods towards ICL-transformers.
2 Related Works
There are three main branches of tools for tabular data classification: classical statistical methods like linear regression, K-nearest neighbors, and support vector machines [17]; tree-based algorithms like XGBoost [6], CatBoost [37], and LightGBM [24]; and neural-network based methods like this paper. There are several papers benchmarking the different methods [12, 43, 15, 32, 50]. Overall, tree-based methods stand at the top, with neural networks ranging from inferior to at best competitive.
Nonetheless, there have been numerous approaches that tackle tabular data classification with neural networks. First, we have the class of neural networks trained from scratch: training starts from random initialized weights and is only trained on the data at hand. Research has focused on architectures [23, 44, 1, 14, 21, 4], embeddings [39, 13, 5], and regularization [42, 22].
In general, methods training from scratch can struggle because tabular datasets can be small. So, researchers have sought ways to use large volumes of tabular data or to change the training objective. Some employ self-supervised learning [26, 49, 54, 2, 46, 45], or closely related transfer learning techniques [33, 27, 53]. Others leverage pretrained LLMs [18, 51] to come to their prediction.
One of those related transfer learning methodologies is in-context learning, a new field sparked by TabPFN [20]. Currently, there is ongoing research on how to scale TabPFN to encompass more observations and features [30, 10, 11]. Amid these developments, we present our work.
3 Preliminaries
In tabular classification, we are interested in predicting targets $y\in{\mathbb{N}}$ given features ${\bm{x}}\in{\mathbb{R}}^{d}$, where $d$ is the number of features. We predict $y$ using an in-context learning (ICL) transformer pretrained on a synthetic dataset. The in-context learning allows the transformer to predict targets based on other observations included in the forward pass. In our work, zero-shot refers to one forward pass through the ICL-transformer without any fine-tuning, while fine-tune refers to one forward pass through the ICL-transformer after fine-tuning. Our work builds on TabPFN [20], so below we explain their dataset generator and their transformer architecture.
3.1 TabPFN Dataset Generator
The TabPFN authors create their own synthetic dataset using Bayesian Neural Networks (BNN) and Structural Causal Models (SCM). To construct a dataset, they first create a BNN or a SCM with random characteristics and with randomly initialized weights. Then they randomly draw an input ${\bm{X}}$ and pass it through the model to generate output ${\bm{y}}$. Their final dataset is given by $({\bm{X}},{\bm{y}})$. See their paper [20] for more details.
In their approach, they emphasize their ability to create realistic datasets, and even call their generator a "prior". They chose SCMs specifically because it can capture real-world causal mechanisms. One other aspect they focus on is simplicity, biasing the generator towards less complex input-output relationships. Additionally, they ensure the inputs ${\bm{X}}$ have natural correlation by correlating the features blockwise, and they vary their feature importance by tuning the magnitude of weights belonging to different features. These methods suggest the authors believe creating realistic datasets is important for achieving good performance.
3.2 Architecture
In our work, we use the architecture from TabPFN, and make no changes to isolate the effect of the dataset generator. This ICL-transformer has as input the features ${\bm{X}}_{support}\in{\mathbb{R}}^{|S|\times d_{f}}$ and targets ${\bm{y}}_{support}\in{\mathbb{N}}^{|S|}$ from support set $S$ and features ${\bm{X}}_{query}\in{\mathbb{R}}^{|Q|\times d_{f}}$ from query set $Q$. The output is a prediction for ${\bm{y}}_{query}\in{\mathbb{R}}^{|Q|}$. The query set $Q$ represents the observations we want to predict, while the support set $S$ are the observations we base our prediction on. This architecture accepts a fixed number of features $d_{f}$, see also the preprocessing discussed in Appendix A.3.
In this transformer, a token with dimension $d_{token}$ represents all features of a single observation. The creation of support tokens $H_{support}\in{\mathbb{R}}^{|S|\times d_{token}}$ and query tokens $H_{query}\in{\mathbb{R}}^{|Q|\times d_{token}}$ are given by equations 1 and 2.
${\bm{H}}_{support}$ $={\bm{X}}_{support}{\bm{W}}_{x}+{\bm{y}}_{support}{\bm{w}}_{y}^{T}$ (1)
${\bm{H}}_{query}$ $={\bm{X}}_{query}{\bm{W}}_{x}$ (2)
Here, we embed the features linearly using weights ${\bm{W}}_{x}\in{\mathbb{R}}^{d_{f}\times d_{token}}$. Input classes ${\bm{y}}_{support}$ are also embedded using a linear layer with weights ${\bm{w}}_{y}\in{\mathbb{R}}^{d_{token}}$, in which ${\bm{y}}_{support}$ is treated as a float. Biases are used but omitted in the equations for conciseness. In Figure 2, $E_{x}$ refers to the multiplication with ${\bm{W}}_{x}$ and $E_{y}$ represents the product with ${\bm{w}}_{y}$.
After the embedding, we push the tokens through a standard transformer architecture with a special attention mask. Support tokens are only able to see other support tokens, and query tokens can only see all support tokens and themselves, with no attention to other query tokens. This attention mask ensures the prediction of an observation does not depend on which other test observations are included in the forward pass. The complete architecture is given in Figure 2.
Figure 2: Base ICL-transformer architecture. On the left, dataset features and targets are separately encoded into tokens. On the right, the targets of the query dataset are used as label. In the middle is the ICL-transformer with the attention mask.
4 Methodology
The full tabular data classification pipeline is given by: synthetic data generation (3.1 and 4.1), data preprocessing (A.3), architectural design (3.2) and fine-tuning (4.2). In this section, we introduce our new dataset generator and our proposed fine-tuning procedure.
4.1 Forest Dataset Generation
Our goal is to create a simple dataset generator that generates datasets with complex patterns to train on, in contrast to the TabPFN [20] generator that aims to create realistic datasets. We base our generator on decision trees, because of their ability to create highly complex decision boundaries with minimal computation cost. The idea is to "overfit" the decision tree to randomly generated features and targets. This fitted decision tree is then used as a data-generating process. See Algorithm 1 for the method and Figure 3 for examples of generated data.
Algorithm 1 Forest Dataset Generation
n_classes, n_features, base_size, dataset_size, tree_depth, categorical_perc
Draw ${\bm{X}}\sim{\mathcal{N}}$(base_size, n_features)
Draw ${\bm{y}}\sim{\mathcal{N}}$(base_size)
Fit a decision tree on (${\bm{X}}$, ${\bm{y}}$) of depth tree_depth.
Draw ${\bm{X}}_{2}\sim{\mathcal{N}}$(dataset_size, n_features)
Convert categorical_perc features of ${\bm{X}}_{2}$ to categorical.
Predict ${\bm{y}}_{2}$ using the decision tree on ${\bm{X}}_{2}$.
Transform ${\bm{y}}_{2}$ using quantile transformation to uniform.
Discretize ${\bm{y}}_{2}$ into n_classes classes.
(${\bm{X}}_{2}$, ${\bm{y}}_{2}$)
Our forest dataset generator allows datasets to vary in number of classes, observations, numerical features, and categorical features. There are two parameters that contribute to the decision boundary complexity. The base size is the number of observations used to fit the decision tree; more observations means more places for the decision tree to split on. The tree depth determines how deep the decision tree will go before exiting the fitting algorithm, with higher depth leading to greater complexity.
For every new synthetically generated dataset, we uniformly draw the hyperparameters from the bounds shown in Table 1. Because we have two hyperparameters that influence the complexity, we decided to keep the base size fixed. In the final step of the algorithm, the targets ${\bm{y}}_{2}$ are discretized by uniformly drawing bucket boundaries between 0.0 and 1.0, and assigning a class to each bucket, creating varying degrees of class imbalance.
Table 1: Hyperparameters for the Forest Dataset Generator
Hyperparameter min max
base size 1024 1024
dataset size 128 1024
tree depth 1 25
number of features 3 100
number of classes 2 10
ratio of categorical features 0.0 1.0
Figure 3: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). The data clouds look unrealistic: decision boundaries are always orthogonal, and there is no feature correlation. Generated with base size 1024, dataset size 1024, maximum tree depth between 1 and 25, two features, and between 2 and 10 number of classes.
4.2 Fine-tuning Procedure
In our work, we introduce fine-tuning to the tabular ICL-transformer. When fine-tuning, we like to draw support and query sets from our training data such that the performance generalizes to the test set. This requires careful consideration of dataset splitting. The benchmark datasets already provide us with a training-validation-test split, of which the validation set is used for hyperparameter search. We require an additional validation set for early stopping, so we split the training with ratio 80/20 in training-small and validation-early-stopping.
Every gradient decent step, we randomly draw a 80/20 support and query split from training-small. For validation, we draw the support set from training-small, and draw the query set from validation-early-stopping. Every sample of the validation-early-stopping is seen exactly once, while support samples are randomly drawn with replacement. We use early stopping based on the validation loss, which is calculated after every step.
The early stopping technique can decide that training zero epochs is optimal, which allows us to fall back on the zero-shot performance in case fine-tuning harms the performance. This is especially important when using very small datasets. At the same time, fine-tuning can make use of the full training dataset, while the number of samples zero-shot can use is limited by the maximum support size that fits on the GPU.
5 Experiments
In our results we consider four pre-trained architectures, each with a zero-shot and a fine-tuned version:
  • TabPFN (original) is the original implementation by the TabPFN [20] authors, fine-tuned by us. The weights are downloaded from their Github.
  • TabPFN (retrained) is trained by us on the TabPFN-dataset.
  • TabForest is trained by us on our forest dataset.
  • TabForestPFN is trained by us on both the TabPFN-dataset and our forest dataset.
Comparing the difference in behavior and performance allows us to learn the effect of the different synthetic datasets. Training and hyperparameter settings are given in appendix A.4, benchmarks used and results obtained are given below.
5.1 Introduction of the Benchmark Datasets
Table 2: WhyTrees Results. Normalized accuracy for mixed and numerical features as shown in Figure 4.
Mixed Numerical
TabPFN (original) Zero-shot 0.530 0.587
TabPFN (original) Fine-tune 0.716 0.738
TabPFN (retrained) Zero-shot 0.207 0.351
TabPFN (retrained) Fine-tune 0.731 0.750
TabForest Zero-shot 0.208 0.350
TabForest Fine-tune 0.829 0.804
TabForestPFN Zero-shot 0.402 0.570
TabForestPFN Fine-tune 0.848 0.835
We show the results of our pre-trained architectures on two benchmarks, tested against publicly available results provided by the authors of the benchmarks. We include all their tested methods and datasets where possible. Appendix A.5 lists all used datasets.
The TabZilla [32] benchmark tests 20 algorithms on 176 classification datasets with sizes ranging from 32 observations to over a million. We selected 94 out of 176 datasets, see appendix A.5. The medium size benchmark which we refer to as WhyTrees [15] consists of 23 classification datasets with 2923 to a maximum of 10,000 observations. The benchmark is split into 7 datasets with only numerical features and 16 datasets with both numerical and categorical features.
Both benchmarks perform random hyperparameter search on their algorithms. TabZilla runs up to 30 times per algorithm and WhyTrees runs a few hundred times, up to 2500 runs. The ICL-transformers run only on default settings because we noticed little gains in performance when changing the fine-tuning hyperparameters.
5.2 Main Results of TabForestPFN
The results on the TabZilla benchmark are shown in Table 3, see Appendix A.6 for alternative presentations. For the WhyTrees benchmark, the comparison of fine-tuned TabForestPFN with the benchmark algorithm is shown in Figure 4, and the comparison with other ICL-transformer variants in Table 2.
Table 3: Main Results on TabZilla. N. Accuracy stands for Normalized accuracy. Rank compares the relative rank of a method compared to all other methods on that dataset. Only the 5 best methods out of 26 are shown. Full table with all methods is presented in Appendix A.6
Models Rank N. Accuracy
min max mean median mean median
TabForestPFN - Fine-tune 1 24 8.0 6.0 0.680 0.663
TabPFN (retrained) - Fine-tune 1 26 8.3 7.8 0.678 0.700
CatBoost 1 22 8.9 7.8 0.676 0.663
XGBoost 1 23 9.2 8.2 0.674 0.671
TabPFN (original) - Fine-tune 1 25 9.3 8.0 0.670 0.677
Figure 4: Main results on the WhyTrees Benchmark. TabForestPFN shows the mean over ten default runs for different fine-tuning seeds, all others use random search over the hyperparameters. See Table 2 for other ICL-transformers.
In these figures and tables, we see that the fine-tuned TabForestPFN outperforms all other methods on TabZilla and matches the performance of XGBoost with a thousand runs on mixed features and a hundred runs on numerical features on WhyTrees. For both benchmarks, we can also see that there is a significant gap between the zero-shot performance of our retrained TabPFN and the original TabPFN. We attribute this phenomenon to the training settings, see Appendix A.4
5.3 Similarity of TabPFN and TabForest
Now that we have established the strong performance of TabForestPFN, we take a closer look at the behavior of TabPFN versus TabForest. In Tables 4 and 6, we see that there is no clear favorite between the fine-tuned versions of TabPFN and TabForest: TabForest is the best on WhyTrees, while TabPFN is favored on TabZilla. Furthermore, TabForestPFN only has a marginal improvement over TabPFN on TabZilla, and similarly only a small performance increase over TabForest on WhyTrees.
In Figure 5 we show the performance of TabPFN, TabForest and TabForestPFN on individual datasets from the WhyTrees benchmark. Remarkably, all three methods often show similar performance across datasets, even though TabPFN and TabForest are pretrained on completely different synthetic data. These results suggest the TabPFN’s performance does not come from the fact that the synthetic dataset is a "prior". Instead the results suggest the performance is a property of an ICL-transformer’s learning process.
Furthermore, we see that there are some differences on some individual datasets, such as the eye movements and, to a lesser extent, the bank-marketing and MiniBooNE datasets. Although we cannot explain why these differences exist, it does prompt us to ensemble the synthetic datasets to create TabForestPFN, which seems to mimic the best method of the two.
Figure 5: Comparison of TabForest, TabPFN and TabForestPFN on 10 datasets of the WhyTrees benchmark. Performance of different ICL-transformers is often similar. All the benchmark algorithms are greyed out. See Appendix A.7 for all WhyTrees datasets and for colored benchmark lines.
5.4 Complexity of ICL-Transformers’ Decision Boundaries
In the previous section, we have seen TabForest and TabPFN show similar performance behavior. Now we take a look at their decision boundaries. Repeating the analysis of the WhyTrees’ authors [15], we use Electricity dataset (OpenML ID: 44156) to predict a target on two features.
To capture the complexity of the decision boundary, we define the complexity score $V$. We split the feature space into a total of $n$ grid cells where each cell has a predicted probability of class one $p_{ij}$ for grid cell indices $(i,j)$. The complexity score $V$ is defined as the sum of absolute values between neighbor cells:
$V=\frac{1}{n}\sum_{ij}|p_{i+1,j}-p_{ij}|+|p_{i-1,j}-p_{i,j}|+|p_{i,j+1}-p_{ij}|+|p_{i,j-1}-p_{ij}|$
The complexity score represents how fast the prediction changes when moving along the grid.
We plot the results in Figure 1. We see that when fine-tuning, both TabPFN and TabForest can create decision boundaries that are more complex than regular MLPs. The complexity of the decision boundaries was one of the characteristics that explained why tree-based methods outperformed neural networks [15]. These results suggest ICL-transformers can also create complexity in the decision boundaries. In Appendix A.10 we also show that the complexity scales with the dimension of the ICL-transformer.
In our intuition, the ICL-transformer learns how to create these decision boundaries during pretraining. Different parameters of the network might be responsible for different parts of the decision boundary. During zero-shot, the ICL-transformer can then create the decision boundary out of these parts. During fine-tuning, the effect is stronger, as different parts come to the surface while irrelevant parts are tuned down.
5.5 Ablation of the Forest Dataset Generator
In Section 4.1 we discussed two ways to influence the complexity of the forest dataset generator: the tree depth and the base size, which is the number of observations to fit the tree algorithm. We expect the performance of TabPFN to increase when the complexity of the forest dataset generator increases.
In Figure 6 we show the results of pretraining different settings of base size and maximum tree depth on the WhyTrees benchmark. The tree depth is set to 1-25 as the base size changes, and the base size is fixed to 1024 as the tree depth changes. When scaling up the base size from 8 to 32 and the tree depth from 1 to 9, we observe that the performance increases, and stabilizes for higher complexities. We provide figures of the data generated with these lower complexity hyperparameters in Appendix A.9 to give an impression. The correlation between performance and complexity supports our claim that learning complex decision boundaries is the driving force behind performance.
Figure 6: Ablation of the base size and maximum tree depth parameters of the Forest Dataset Generator. Figure shows normalized test accuracy of TabForest on the WhyTrees benchmark. A smaller model is used (dim: 512 -> 256, layers: 12 -> 8).
5.6 Superiority of Fine-tuning over Zero-shot
In the main results, we have seen that fine-tuning performs better than zero-shot. We look at this comparison in more detail. Figure 7(a) presents the performance of TabForestPFN on individual datasets from TabZilla. We can see clearly that fine-tuning strongly outperforms zero-shot when there are more than 10,000 observations. However, fine-tuning struggles on the two smallest datasets, which have four and six validation samples. Overall, fine-tuning outperforms zero-shot on 66% of the datasets. This percentage decreases to 54% for datasets smaller than a 1000 observations, and increases to 84% to datasets larger than a 1000 observations.
Figure 7(b) illustrates the effect of context length on the performance of TabForestPFN on the zero-shot and the fine-tuning task. We see that a higher support size is always better, which is why we set the support size in our paper to 8192, even though we only pretrained on a maximum size of 1024 observations. In conclusion, fine-tuning on the largest possible support size appears to be the most effective approach for ICL-transformers.
(a) Differences in normalized accuracy of individual datasets from TabZilla. The color red means fine-tuning is the best. The darkest red represents at least 0.20 normalized score points improvement, and dark blue at least 0.20 normalized accuracy points degradation.
(b) Effect of different support sizes on the WhyTrees benchmark. Both fine-tuning and zero-shot performance improves with context size.
Figure 7: Evaluation of TabForestPFN
6 Conclusion
The introduction of TabPFN [20] has opened up a new field of in-context learning (ICL)-transformers for tabular data classification. Our research has demonstrated that these ICL-transformers learn to create complex decision boundaries. Furthermore, by fine-tuning and ensembling the pretraining dataset, we achieved performance levels that surpass XGBoost [6] and CatBoost [37] on third-party benchmarks [15, 32]. Although tree based-methods can continue to gain performance through hyperparameter optimization, we expect ICL-transformers can match their gains by scaling up the pretraining process.
Despite these advancements, there are still obstacles to overcome for ICL-transformers to fully replace tree-based methods in the realm of tabular data. One major challenge is the performance limitation of ICL-transformers due to GPU memory constraints. Our work uses fine-tuning as a solution to this problem, but it would be valuable to compare this approach to other concurrent research such as prompt tuning [11] and context distillation [30]. Moreover, our research solely focused on classification, although we expect a simple switch from cross-entropy loss to mean-squared-error loss would suffice to tackle regression tasks. Another area that requires exploration is the setting with an exceptionally high number of features [7], where the performance of ICL-transformers is unknown [31]. Lastly, tree-based methods have an edge in explaining the feature importance, and research needs to be done if ICL-transformers can achieve similar interpretability [40]. Overcoming these challenges will cement the ICL-transformer as the clear successor of tree-based methods.
Appendix A Appendix
A.1 Ethics and Social Impact
Improving tabular data classification can provide major benefits to society. From medical to physics applications, better performance can save lives and money. There are, however, also more nefarious applications of tabular data classification, such as fraud risk detections based on ethnics or nationality; population analysis for micro-targeting political ad campaigns; and insurance premium discrimination based on underlying medical conditions.
Our models cannot detect the purpose for which the model is used. In contrast to large language models, our tabular data models take numerical data as input. Ethnicity, gender, or other sensitive information is represented by a class number. This means our models cannot recover the meaning behind the numbers.
A big benefit of this ’numerical anonymization’ is privacy. In our paper, we use synthetic data, which is completely privacy risk free. But even when pretrained on real data, recovering the original data can be extremely hard due to the lack of labels and contextual information.
In light of the above, we decide to publish the model and open access to anyone. We do not know an effective way to create any form of safeguards against misuse, and we would welcome any advice from the research community that addresses this issue.
A.2 Code Availability
Code is available at https://github.com/FelixdenBreejen/TabForestPFN
The code includes everything: downloading datasets, preprocessing result benchmarks, training the ICL-transformers, pretrained weights on Google Drive, notebooks with analysis, and an easy example to get you started with applying the TabForestPFN to your dataset.
The code is built upon the works of Grinsztajn et al., Hollmann et al. and McElfresh et al.. Although this resulted in a Frankenstein monster of a codebase, we have made great efforts to rewrite most of their code to integrate well. The code assumes a single server with access to 1 or more GPUs, with DistributedDataParallel used during pretraining and multiprocessing over different GPUs used during inference.
As currently there is no good tabular code base out there, we recommend anyone that is interested in doing research in tabular data to take a look at ours.
A.3 Data Preprocessing
Before data is put into the neural network, the data is preprocessed. We use the exact same routine for both synthetic data and real-world data to ensure minimal differences in distribution and summary statistics of the input to the transformer. Algorithm 2 presents the procedure for preprocessing.
Algorithm 2 Data Preprocessing
1:${\bm{X}}_{raw}$, ${\bm{y}}_{raw}$
2:Impute NaN features with column mean.
3:Remove features with one unique value.
4:Select a subset of a hundred features.
5:Transform all features to normal using quantile transformation.
6:Normalize data to unit mean and variance.
7:Scale data based on number of features.
8:Pad the features to $d_{f}$ features by adding zeros.
9:if Pretraining then
10:     Shuffle the order of the features and classes.
11:end if
12:${\bm{X}}_{preprocessed}$, ${\bm{y}}_{preprocessed}$
Because we fixed the input size of the neural network to $d_{f}=100$ features, we first select a subset of a $d_{f}$ features using scikit-learn’s SelectKBest [36]. If there are less than $d_{f}$ features, we add zeros to ensure exactly $d_{f}$ features. Following TabPFN, we scale by multiplying by $100/d_{f*}$, where $d_{f*}$ is the number of features after selecting a subset. To be robust to skewness and outliers, we transform the data using scikit-learn’s QuantileTransformer to follow a normal distribution. We make no distinction between numerical and categorical values in all our preprocessing.
In comparison to TabPFN, our data preprocessing follows roughly the same scheme. One change is the use of a quantile transformer, while they use standard input or a power transformer. We consider this a preference, both seem to work fine.
Furthermore, the TabPFN authors like to ensemble outputs of the TabPFN architecture, varying the transformation function between standard input and power transformer. In our paper, we use no ensembling at all.
A.4 Training Settings
All ICL-transformer architectures, including the original TabPFN, use the same model. The model consists of 12 layers, 4 attention heads, a hidden dimension of 512, and 10 classes as output dimension. Pretraining uses batch size 64, learning rate 1e-4, weight decay 0.00, AdamW optimizer with betas (0.9, 0.95), cosine scheduling, and maximum global gradient norm 1.0. Fine-tuning is performed under batch size 1, learning rate 1e-5, weight decay 0.00, no scheduling, with early stopping, and a maximum of 300 steps. To fit the model on an RTX 3090 with 24GB during fine-tuning, we set the maximum support size to 8192 samples and the maximum query size to 1024. Pre-training uses data generated with maximum support size 1024 and maximum query size 128. We choose these settings because the maximum support size affects performance, but the maximum query size only affects inference speed.
On these settings, we train TabPFN and TabForest for 300,000 steps, while TabForestPFN trains for 600,000 steps. Both TabPFN and TabForest can likely train longer; however, we experienced high instability in performance after 300,000 steps depending on the seed. Sometimes, the fine-tuning ability of the model would completely collapse, and sometimes the performance would slowly overfit. We also chose 300,000 steps because we did not see improvement in fine-tuning performance in the steps afterward. We chose to use 600,000 steps for the combined TabForestPFN, because we believed double the data would allow us to use double the steps, and we experienced no instability on this setting. Running 600,000 steps on one RTX 3090 takes 24 GPU-days.
Choosing 300,000 steps does have a downside. In the results in section 5.2, we saw the retrained TabPFN and the TabForest have poor zero-shot performance. The number of steps therefore is a trade-off: a small number of steps is stable and has high fine-tuning performance, while a large number of steps has good zero-shot performance. Because in this paper we are interested in maximum performance, we prioritize the fine-tuned version on 300,000 steps.
A.5 Benchmark Metadata
Of both the TabZilla and the WhyTrees benchmark, we show the OpenML[47] datasets we use as well as their characteristics. See Table LABEL:table:metadata_tabzilla and Table LABEL:table:metadata_whytrees. The TabZilla table presents the 94 datasets picked out of the 176 total datasets.
From the original 176 Tabzilla datasets, we excluded every dataset that does not have at least one completed run on default settings for every model, which brings the value to 99. Additionally, we exclude four datasets because they have more than 10 classes. The preprocessing code of one other dataset did not run without errors, and so is removed as well. The TabZilla authors did experiment with running TabPFN, but only on 62 datasets with a maximum support size of 3000 samples, so we redo their experiment.
Table 4: Metadata of the TabZilla Benchmark. Splits refers to the number of cross validation splits.
OpenML Observations Features Splits Classes
ID Name All Train Valid Test
3 kr-vs-kp 3196 2556 320 320 36 10 2
4 labor 57 45 6 6 16 10 2
9 autos 205 163 21 21 25 10 6
10 lymph 148 118 15 15 18 10 4
11 balance-scale 625 499 63 63 4 10 3
12 mfeat-factors 2000 1600 200 200 216 10 10
14 mfeat-fourier 2000 1600 200 200 76 10 10
15 breast-w 699 559 70 70 9 10 2
16 mfeat-karhunen 2000 1600 200 200 64 10 10
18 mfeat-morpholog… 2000 1600 200 200 6 10 10
23 cmc 1473 1177 148 148 9 10 3
25 colic 368 294 37 37 26 10 2
27 colic 368 294 37 37 22 10 2
29 credit-approval 690 552 69 69 15 10 2
30 page-blocks 5473 4377 548 548 10 10 5
35 dermatology 366 292 37 37 34 10 6
37 diabetes 768 614 77 77 8 10 2
39 sonar 208 166 21 21 60 10 2
40 glass 214 170 22 22 9 10 6
43 spambase 4601 3680 460 461 57 10 2
45 splice 3190 2552 319 319 60 10 3
47 tae 151 120 15 16 5 10 3
48 heart-c 303 241 31 31 13 10 2
49 tic-tac-toe 958 766 96 96 9 10 2
50 heart-h 294 234 30 30 13 10 2
53 vehicle 846 676 85 85 18 10 4
59 iris 150 120 15 15 4 10 3
2074 satimage 6430 5144 643 643 36 10 6
2079 eucalyptus 736 588 74 74 19 10 5
2867 anneal 898 718 90 90 38 10 5
3485 scene 2407 1925 241 241 299 10 2
3512 synthetic_contr… 600 480 60 60 60 10 6
3540 analcatdata_box… 120 96 12 12 3 10 2
3543 irish 500 400 50 50 5 10 2
3549 analcatdata_aut… 841 672 84 85 70 10 4
3560 analcatdata_dmf… 797 637 80 80 4 10 6
3561 profb 672 536 68 68 9 10 2
3602 visualizing_env… 111 88 11 12 3 10 2
3620 fri_c0_100_5 100 80 10 10 5 10 2
3647 rabe_266 120 96 12 12 2 10 2
3711 elevators 16599 13279 1660 1660 18 10 2
3731 visualizing_liv… 130 104 13 13 2 10 2
3739 analcatdata_chl… 100 80 10 10 3 10 2
3748 transplant 131 104 13 14 3 10 2
3779 fri_c3_100_5 100 80 10 10 5 10 2
3797 socmob 1156 924 116 116 5 10 2
3896 ada_agnostic 4562 3648 457 457 48 10 2
3902 pc4 1458 1166 146 146 37 10 2
3903 pc3 1563 1249 157 157 37 10 2
3904 jm1 10885 8707 1089 1089 21 10 2
3913 kc2 522 416 53 53 21 10 2
3917 kc1 2109 1687 211 211 21 10 2
3918 pc1 1109 887 111 111 21 10 2
3953 adult-census 32561 26048 3256 3257 14 10 2
9946 wdbc 569 455 57 57 30 10 2
9952 phoneme 5404 4322 541 541 5 10 2
9957 qsar-biodeg 1055 843 106 106 41 10 2
9960 wall-robot-navi… 5456 4364 546 546 24 10 4
9964 semeion 1593 1273 160 160 256 10 10
9971 ilpd 583 465 59 59 10 10 2
9978 ozone-level-8hr 2534 2026 254 254 72 10 2
9984 fertility 100 80 10 10 9 10 2
10089 acute-inflammat… 120 96 12 12 6 10 2
10093 banknote-authen… 1372 1096 138 138 4 10 2
10101 blood-transfusi… 748 598 75 75 4 10 2
14952 PhishingWebsite… 11055 8843 1106 1106 30 10 2
14954 cylinder-bands 540 432 54 54 37 10 2
14965 bank-marketing 45211 36168 4521 4522 16 10 2
14967 cjs 2796 2236 280 280 33 10 6
125920 dresses-sales 500 400 50 50 12 10 2
125921 LED-display-dom… 500 400 50 50 7 10 10
145793 yeast 1269 1015 127 127 8 10 4
145799 breast-cancer 286 228 29 29 9 10 2
145836 blood-transfusi… 748 598 75 75 4 10 2
145847 hill-valley 1212 968 122 122 100 10 2
145977 ecoli 336 268 34 34 7 10 8
145984 ionosphere 351 280 35 36 34 10 2
146024 lung-cancer 32 24 4 4 56 10 3
146063 hayes-roth 160 128 16 16 4 10 3
146065 monks-problems-… 601 480 60 61 6 10 2
146192 car-evaluation 1728 1382 173 173 21 10 4
146210 postoperative-p… 88 70 9 9 8 10 2
146607 SpeedDating 8378 6702 838 838 120 10 2
146800 MiceProtein 1080 864 108 108 77 10 8
146817 steel-plates-fa… 1941 1552 194 195 27 10 7
146818 Australian 690 552 69 69 14 10 2
146820 wilt 4839 3871 484 484 5 10 2
146821 car 1728 1382 173 173 6 10 4
167140 dna 3186 2548 319 319 180 10 3
167141 churn 5000 4000 500 500 20 10 2
167211 Satellite 5100 4080 510 510 36 10 2
168911 jasmine 2984 2386 299 299 144 10 2
190408 Click_predictio… 39948 31958 3995 3995 11 10 2
360948 libras 360 288 36 36 104 10 10
Table 5: Metadata of the WhyTrees Benchmark. Splits refers to the number of cross validation splits.
OpenML Observations Features Splits Classes
ID Name All Train Valid Test
44089 credit 16714 10000 2014 4700 10 2 2
44120 electricity 38474 10000 8542 19932 7 1 2
44121 covertype 566602 10000 50000 50000 10 1 2
44122 pol 10082 7057 907 2118 26 3 2
44123 house_16H 13488 9441 1214 2833 16 3 2
44125 MagicTelescope 13376 9363 1203 2810 10 3 2
44126 bank-marketing 10578 7404 952 $2$ 7 3 2
44128 MiniBooNE 72998 10000 18899 44099 50 1 2
44129 Higgs 940160 10000 50000 50000 24 1 2
44130 eye_movements 7608 5325 684 1599 20 3 2
44156 electricity 38474 10000 8542 19932 8 1 2
44157 eye_movements 7608 5325 684 1599 23 3 2
44159 covertype 423680 10000 50000 50000 54 1 2
45019 Bioresponse 3434 2403 309 722 419 5 2
45020 default-of-cred… 13272 9290 1194 2788 20 3 2
45021 jannis 57580 10000 14274 33306 54 1 2
45022 Diabetes130US 71090 10000 18327 42763 7 1 2
45026 heloc 10000 7000 900 2100 22 3 2
45028 california 20634 10000 3190 7444 8 1 2
45035 albert 58252 10000 14475 33777 31 1 2
45036 default-of-cred… 13272 9290 1194 2788 21 3 2
45038 road-safety 111762 10000 30528 50000 32 1 2
45039 compas-two-year… 4966 3476 447 1043 11 3 2
A.6 TabZilla Further Results
Table 6: Main Results on TabZilla. N. Accuracy stands for Normalized accuracy. Rank compares the relative rank of a method compared to all other methods on that dataset.
Models Rank N. Accuracy
min max mean median mean median
TabForestPFN - Fine-tune 1 24 8.0 6.0 0.680 0.663
TabPFN (retrained) - Fine-tune 1 26 8.3 7.8 0.678 0.700
CatBoost 1 22 8.9 7.8 0.676 0.663
XGBoost 1 23 9.2 8.2 0.674 0.671
TabPFN (original) - Fine-tune 1 25 9.3 8.0 0.670 0.677
TabForestPFN - Zero-shot 1 26 10.6 10.0 0.633 0.650
TabForest - Fine-tune 1 24 10.6 9.5 0.650 0.667
TabPFN (original) - Zero-shot 1 25 10.9 10.2 0.632 0.606
LightGBM 1 26 11.1 11.0 0.646 0.653
RandomForest 1 25 11.3 11.0 0.636 0.637
Resnet 1 26 11.9 10.0 0.602 0.613
NODE 1 26 12.1 12.0 0.611 0.596
SAINT 1 26 12.2 12.8 0.600 0.614
SVM 1 25 12.5 13.2 0.589 0.573
FT-Transformer 1 23 12.6 12.5 0.599 0.601
DANet 3 25 14.5 15.0 0.596 0.608
MLP-rtdl 1 26 15.8 17.8 0.514 0.527
TabForest - Zero-shot 3 25 15.9 17.5 0.571 0.583
TabPFN (retrained) - Zero-shot 3 25 15.9 17.5 0.571 0.583
STG 1 26 16.1 17.2 0.495 0.472
LinearRegression 1 26 17.5 19.5 0.459 0.431
MLP 1 26 17.7 20.0 0.475 0.417
TabNet 2 26 18.0 19.0 0.503 0.490
DecisionTree 1 26 18.7 21.0 0.413 0.372
KNN 2 26 19.4 21.5 0.399 0.372
VIME 1 26 21.8 23.8 0.295 0.221
In the TabZilla main results Table 3, we have shown the five highest performing methods. The full results are shown in Table 6, which includes all ICL-transformer variants and methods implemented by the TabZilla authors. Because the rank is calculated over all included methods, which ICL-transformer variants we include might change the results. Therefore, we check if the results are the same if we use calculate the rankings one ICL-transformer at the time.
Table 7 shows the results of only the fine-tuned TabForestPFN versus the rest of the benchmark. We do this for every ICL-transformer and aggregregate the results in Table 8. All results are qualitatively the same as in Table 3.
Table 7: Main Results on TabZilla. N. Accuracy stands for Normalized accuracy. Rank compares the relative rank of a method compared to all other methods on that dataset.
Models Rank N. Accuracy
min max mean median mean median
TabForestPFN - Fine-tune 1 18 5.8 4.2 0.680 0.663
CatBoost 1 15 6.2 5.2 0.676 0.663
XGBoost 1 16 6.3 5.0 0.674 0.671
LightGBM 1 19 7.6 6.0 0.646 0.653
RandomForest 1 18 7.9 8.0 0.636 0.637
NODE 1 19 8.4 8.0 0.611 0.596
Resnet 1 19 8.4 8.0 0.602 0.613
SAINT 1 19 8.5 7.2 0.600 0.614
SVM 1 18 8.6 8.0 0.589 0.573
FT-Transformer 1 16 8.8 8.5 0.599 0.601
DANet 2 19 10.0 10.0 0.596 0.608
MLP-rtdl 1 19 11.4 12.0 0.514 0.527
STG 1 19 11.5 12.0 0.495 0.472
LinearRegression 1 19 12.3 13.8 0.459 0.431
MLP 1 19 12.6 14.0 0.475 0.417
TabNet 1 19 12.8 13.8 0.503 0.490
DecisionTree 1 19 13.3 14.0 0.413 0.372
KNN 2 19 13.9 15.0 0.399 0.372
VIME 1 19 15.7 17.0 0.295 0.221
Table 8: Main Results on TabZilla. N. Accuracy stands for Normalized accuracy. Rank compares the relative rank of a method compared to all other methods on that dataset. This table displays individual results: Table 7 is run individually for all ICL-transformer variants, and the row of the ICL-transformer is copy-pasted here.
Models Rank N. Accuracy
min max mean median mean median
TabPFN (original) - Zero-shot 1 19 7.4 6.0 0.632 0.606
TabPFN (original) - Fine-tune 1 19 6.4 5.0 0.670 0.677
TabPFN (retrained) - Zero-shot 2 19 10.7 11.8 0.571 0.583
TabPFN (retrained) - Fine-tune 1 19 5.9 5.0 0.678 0.700
TabForest - Zero-shot 2 19 10.7 11.8 0.571 0.583
TabForest - Fine-tune 1 19 7.4 7.0 0.650 0.667
TabForestPFN - Zero-shot 1 19 7.2 6.0 0.633 0.650
TabForestPFN - Fine-tune 1 18 5.8 4.2 0.680 0.663
A.7 WhyTrees Further Results
The main results in Figure 4 report the normalized accuracy aggregated over all datasets. In Figure 8 we show the comparison between ICL-transformers on all 23 datasets. In Figure 9 we show the colored results on each individual dataset where we compare fine-tuned TabForestPFN with all the benchmark methods.
Figure 8: Comparison of TabForest, TabPFN and TabForestPFN on the WhyTrees benchmark. Performance of different ICL-transformers is often similar. All the benchmark algorithms are greyed out. See Appendix A.7 for colored benchmark lines.
s
Figure 9: Individual dataset results on the WhyTrees Benchmark. TabForestPFN shows the mean of ten default runs, all others use random search over the hyperparameters. See Table 2 for other ICL-transformers.
A.8 One-by-One Comparisons
In Figure 10 we plot one-to-one comparisons of fine-tuned TabForestPFN versus CatBoost, TabForest and TabPFN. We see no clear correlations in the other comparisons between performance difference and model.
Figure 10: Differences in normalized accuracy of individual datasets from TabZilla. The color red means the left-mentioned method is the best, blue for the right-mentioned method. The darkest red represents at least 0.20 normalized score points improvement, and dark blue at least 0.20 normalized accuracy points degradation.
A.9 Synthetic Data with Lower Complexity
In the ablation we have seen that even with a forest dataset generator with lower complexity parameters, we still have similar performance. To give an idea of how complex the data is, here we showcase the generated data. Figure 11 displays generated data with base size 32, and Figure 12 displays generated data with maximum tree depth 9.
Figure 11: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). Generated with base size 32, dataset size 1024, tree depth between 1 and 25, two features, and between 2 and 10 number of classes. See also Figures 3 and 12.
Figure 12: Generated forest data. Every box is a generated dataset with its own classes (color) and features (axes). Generated with base size 1024, dataset size 1024, tree depth between 1 and 9, two features, and between 2 and 10 number of classes. See also Figures 3 and 11.
A.10 Decision Boundaries for Scaled TabForest
In the complexity boundary results, we saw that TabForest has more complex decision boundaries when fine-tuned. Here we show that the complexity of the decision boundary is correlated with the dimension size of the transformer. See Figure 13 for the results. We see that the decision boundaries become more complex when the dimension of the ICL-transformer increases.
Figure 13: Comparison of decision boundaries for the electricity dataset. Axis represent features, color is predicted class probability, dots are test observations. When increasing the dimension size of the transformer, complexity increases.